Find the value of the finite population correction factor for n = 12 and N = 1,000.
1
Expert's answer
2021-08-31T02:49:42-0400
"FPC=\sqrt{\dfrac{N-n}{N-1}}=\sqrt{\dfrac{1000-12}{1000-1}}\approx0.994479"